# MoMa: A Modular Deep Learning Framework for Material Property Prediction

### v2

@@ Abstract
Deep learning methods for material property prediction have been widely explored to advance materials discovery. However, the prevailing pre-train then fine-tune paradigm often fails to address the inherent diversity and disparity of material tasks. To overcome these challenges, we introduce MoMa, a Modular framework for Materials that first trains specialized modules across a wide range of tasks and then adaptively composes synergistic modules tailored to each downstream scenario. Evaluation across 17 datasets demonstrates the superiority of MoMa, with a substantial 14% average improvement over the strongest baseline. Few-shot and continual learning experiments further highlight MoMa’s potential for real-world applications. Pioneering a new paradigm of modular material learning, MoMa will be open-sourced to foster broader community collaboration.
modular deep learning, material property prediction, materials science
@@ 1 Introduction
Accurate and efficient material property prediction is critical for accelerating materials discovery. Key properties such as formation energy and band gap are fundamental in identifying stable materials and functional semiconductors (Riebesell et al., 2023; Masood et al., 2023). While traditional approaches such as density functional theory (DFT) offer high precision (Jain et al., 2016), their prohibitive computational cost limits their practicality for large-scale screening (Fiedler et al., 2022; Lan et al., 2023).
@@
Recently, deep learning methods have been developed to expedite traditional approaches (Xie & Grossman, 2018; Griesemer et al., 2023). Pre-trained force field models, in particular, have shown remarkable success in generalizing to a wide spectrum of material property prediction tasks (Yang et al., 2024b; Barroso-Luque et al., 2024; Shoghi et al., 2023), outperforming specialized models trained from scratch. These models are typically pre-trained on the potential energy surface (PES) data of materials and then fine-tuned for the target downstream task.
Figure 1: Illustration of the diversity of material properties (top) and systems (bottom). Note that material tasks are also disparate, with different laws governing the diverse properties and systems. These characteristics pose challenges for pre-training material property prediction models.
@@
Despite these advances, we identify two key challenges that undermine the effectiveness of current pre-training strategies for material property prediction: diversity and disparity.
@@
First, material tasks exhibit significant diversity (Figure 1), which current pre-trained models fail to adequately cover. Existing models trained on PES-derived properties (e.g., force, energy, and stress) mostly focus on crystalline materials (Yang et al., 2024b; Barroso-Luque et al., 2024). However, material tasks span a wide variety of systems (e.g., crystals, organic molecules) and properties (e.g., thermal stability, electronic behavior, mechanical strength), making it difficult for methods trained on a limited set of data to generalize across the full spectrum of tasks.
@@
Second, the disparate nature of material tasks presents huge obstacles for jointly pre-training a broad span of tasks. Material systems vary significantly in terms of bonding, atomic composition, and structural periodicity, while their properties are governed by distinct physical laws. For example, mechanical strength in metals is primarily influenced by atomic bonding and crystal structure, whereas electronic properties like conductivity are determined by the material’s electronic structure and quantum mechanics. Consequently, training a single model across a wide range of tasks (Shoghi et al., 2023) may lead to knowledge conflicts, hindering the model’s ability to effectively adapt to downstream scenarios.
@@
In this paper, we propose MoMa, a Modular deep learning framework for Material property prediction, to address the diversity and disparity challenge. To accommodate the diversity of material tasks, MoMa first trains on a multitude of high-resource property prediction datasets, centralizing them into transferrable modules. Furthermore, MoMa incorporates an adaptive composition algorithm that customizes support for diverse downstream scenarios. Recognizing the disparity among material tasks, MoMa encapsulates each task within a specialized module, eliminating task interference of joint training. In adapting MoMa to specific downstream tasks, its composition strategy adaptively integrates only the most synergistic modules, mitigating knowledge conflicts and promoting positive transfer.
@@
Specifically, MoMa comprises two major stages: (1) Module Training & Centralization. Drawing inspiration from modular deep learning (Pfeiffer et al., 2023), MoMa trains dedicated modules for a broad range of material tasks, offering two versions: a full module for superior performance and a memory-efficient adapter module. These trained modules are centralized in MoMa Hub, a repository designed to facilitate knowledge reuse while preserving proprietary data for privacy-aware material learning. (2) Adaptive Module Composition (AMC). MoMa introduces the data-driven AMC algorithm that composes synergetic modules from MoMa Hub. AMC first estimates the performance of each module on the target task in a training-free manner, then heuristically optimizes their weighted combination. The resulting composed module is then fine-tuned for improved adaptation to the downstream task. Together, the two stages deliver a modular solution that enables MoMa to account for the diversity and disparity of material knowledge.
@@
Empirical results across 17 downstream tasks showcase the superiority of MoMa, outperforming all baselines in 16/17 tasks, with an average improvement of 14% compared to the second-best baseline. In few-shot settings, which are common in materials science, MoMa achieves even larger performance gains to the conventional pre-train then fine-tune paradigm. Additionally, we show that MoMa can expand its capability in continual learning settings by incorporating molecular tasks into MoMa Hub. The trained modules in MoMa Hub will be open-sourced, and we envision MoMa becoming a pivotal platform for the modularization and distribution of materials knowledge, fostering deeper community engagement to accelerate materials discovery.
@@ 2 Proposed Framework: MoMa
MoMa is a simple modular framework targeting the diversity and disparity of material tasks. The predominant pre-train then fine-tune strategy can only leverage a limited range of interrelated source tasks or indiscriminately consolidate conflicting knowledge into one model, resulting in suboptimal downstream performance. In contrast, the modular design of MoMa allows for the flexible and scalable integration of diverse material knowledge modules, and the effective and tailored adaptation to material property prediction tasks. Figure 2 illustrates this comparison.
Figure 2: A comparison between the pre-train fine-tune paradigm and MoMa’s modular framework. (left): The prevailing scheme involves pre-training on force field data (with supervised prediction on energy, force, and stress), and then transfer to downstream tasks. (right): The modular learning scheme in MoMa trains and stores a broad spectrum of material tasks as modules, and adaptively composes them given a new material property prediction task.
2.1 Overview
MoMa involves two major stages: (1) training and centralizing modules into MoMa Hub; (2) adaptively composing these modules to support downstream material tasks.
In the first stage (Section 2.2), we encompass a wide range of material properties and systems into MoMa Hub. This accommodates the diversity of material tasks and addresses the task disparity by training specialized modules for each.
In the second stage (Section 2.3), we devise the Adaptive Module Composition algorithm. Given the downstream material task, the algorithm heuristically optimizes the optimal combination of module weights for MoMa Hub and composes a customized module based on the weights, which is subsequently fine-tuned on the task for better adaptation. Respecting the diverse and disparate nature of material tasks, our adaptive approach automatically discovers synergistic modules and excludes conflicting combinations by the data-driven assignment of module weights.
A visual overview of MoMa is provided in Figure 3.
Figure 3: The MoMa framework. (a) During the Module Training & Centralization stage (Section 2.2), MoMa trains full and adapter modules for a wide spectrum of material tasks, constituting the MoMa Hub; (b) The Adaptive Module Composition (AMC) & Fine-tuning stage (Section 2.3) leverages the modules in MoMa Hub to compose a tailored module for each downstream task. The AMC algorithm comprises three steps: 1. module prediction estimation (with $k$NN); 2. module weight optimization; 3. module composition. The composed module is further fine-tuned on the task for better adaptation.
@@ 2.2 Module Training & Centralization
We provide two parametrizations for the MoMa modules: the full module and the adapter module. For the full module, we directly treat each fully fine-tuned backbone as a module. The adapter module serves as a parameter-efficient variant where adapter layers (Houlsby et al., 2019) are inserted between each layer of the pre-trained backbone. The adapters are updated and the rest of the backbone is frozen. All of the adapters for each task are treated as one module. This implementation trade-offs the downstream performance in favor of a significantly lower GPU memory cost during training, which is particularly beneficial when the computational resources are constrained. When the training converges, we store the module parameters into a centralized repository $\mathcal{H}$ termed MoMa Hub, formally:
=$\mathcal{H}=\{g_{1},g_{2},\dots,g_{N}\},\quad g_{i}=\begin{cases}\theta_{f}^{i}&\text{(full module)}\\ \Delta_{f}^{i}&\text{(adapter module)}\end{cases}$
where $\theta_{f}^{i}$ and $\Delta_{f}^{i}$ denote the full and adapter module parameters related to the $i^{\text{th}}$ task and encoder $f$.
@@
To support a wide array of downstream tasks, MoMa Hub needs to include modules trained on diverse material systems and properties. Currently, MoMa Hub encompasses 18 material property prediction tasks selected from the Matminer datasets (Ward et al., 2018) with over 10000 data points. These tasks span across a large range of material properties, including thermal properties (e.g. formation energy), electronic properties (e.g. band gap), mechanical properties (e.g. shear modulus), etc. For more details, please refer to Table 3. To showcase the effect of scaling data diversity, we present the continual learning results in Section 3.5 after further incorporating molecular property prediction tasks into MoMa Hub. Note that MoMa is designed to be task-agnostic and may readily support a larger spectrum of tasks in the future.
An important benefit of the modular design of MoMa Hub is that it preserves proprietary data, which is prevalent in the field of materials, enabling privacy-aware contribution of new modules. Therefore, MoMa could serve as an open platform for the modularization of materials knowledge, which also facilitates downstream adaptation through a novel composition mechanism, as discussed in the following section.
@@ 2.3 Adaptive Module Composition & Fine-tuning
Given a labeled material property prediction dataset $\mathcal{D}$ with $m$ instances: $\mathcal{D}=\{(x_{1},y_{1}),(x_{2},y_{2}),\ldots,(x_{m},y_{m})\}$, the second stage of MoMa customizes a task-specific model using the modules in MoMa Hub.
To achieve this, we devise the Adaptive Module Composition (AMC) algorithm. We highlight its key desiderata:
  • Selective: Material tasks are inherently disparate. Hence only the most relevant modules shall be selected to avoid the negative interference of materials knowledge and encourage positive transfer to downstream tasks.
  • Data-driven: As the diversity of tasks in MoMa Hub expands, it is impossible to rely solely on human expertise for module selection. A data-driven approach is required to mine the implicit relationships between the MoMa Hub modules and downstream tasks.
  • Efficient: Enumerating all combinations of modules is impractical. Efficient algorithms shall be developed to return the optimal module composition using a reasonable amount of computational resources.
To meet these requirements, AMC is designed as a fast heuristic algorithm that first estimates the prediction of each module on the downstream task, then optimizes the module weights, and finally composes the selected modules to form the task-specific module. We now elaborate on the details of AMC, with its formal formulation in Algorithm 1.
Module Prediction Estimation
We begin by estimating the predictive performance of each module in MoMa Hub $\mathcal{H}$ on the downstream task $\mathcal{D}$. More accurate predictions indicate stronger relevance to the task and intuitively warrant higher weights in the composition.
For each module $g_{j}$ in $\mathcal{H}$, we first take it to encode each input materials in the train set of task $\mathcal{D}$ into a set of representation $\mathcal{X}^{j}=\{\mathbf{x}_{1}^{j},\mathbf{x}_{2}^{j},\ldots,\mathbf{x}_{m}^{j}\}$ in which $\mathbf{x}_{i}^{j}=g_{j}({x}_{i})$. Then we obtain the estimated prediction of $g_{j}$ on $\mathcal{D}$ using a leave-one-out label propagation approach (Iscen et al., 2019). Specifically, we iteratively select one sample $\mathbf{x}_{i}^{j}$ from $\mathcal{X}^{j}$ and get the predicted label $\hat{y_{i}}^{j}$ by calculating the weighted sum of its $K$ nearest neighbors’ labels within $\mathcal{X}^{j}$:
$\hat{y_{i}}^{j}=\sum_{k=1}^{K}\frac{f_{d}(\mathbf{x}_{i}^{j},\mathbf{x}_{k}^{j})}{Z}y_{k},$ (1)
@@
where $\mathbf{x}_{k}^{j}$ denotes the $k$-th nearest neighbors of $\mathbf{x}_{i}^{j}$. The distance function $f_{d}$ for calculating $k$NN is the exponential of cosine similarity between each pair of $\mathbf{x}_{i}^{j}$ and $\mathbf{x}_{k}^{j}$. $Z=\sum_{k=1}^{K}f_{d}(\mathbf{x}_{i}^{j},\mathbf{x}_{k}^{j})$ is the normalizing term.
While other predictors are viable, we choose $k$NN due to its good trade-off in efficiency and accuracy. Also, its training-free nature enhances its flexibility in real-world scenarios, where the downstream data may be subject to updates.
Module Weight Optimization
After estimating each module’s prediction, we now have to select the optimal combination of modules tailored for the downstream task $\mathcal{D}$. To achieve this, the most straightforward approach is to compare the prediction error obtained after fine-tuning each combination of modules. However, this is infeasible due to the combinatorial explosion. Therefore, we reformulate the task as an optimization problem, using the prediction error before fine-tuning as a proxy metric (later referred to as proxy error). By optimizing the proxy error, we could obtain the optimal combination of weights.
Specifically, inspired by ensemble learning (Zhou et al., 2002; Zhou, 2016), we assign a weight $w_{j}$ for each module $g_{j}$ and calculate the output of the ensemble: $\sum_{j=1}^{NT}w_{j}\hat{y_{i}}^{j}$. We then estimate the proxy error on the train set of $\mathcal{D}$ for this weighted ensemble:
$E_{\mathcal{D}}=\frac{1}{m}\sum_{i=1}^{m}(\sum_{j=1}^{N}w_{j}\hat{y}_{i}^{j}-y_{i})^{2}$ (2)
To minimize the proxy error $E_{\mathcal{D}}$, we then utilize the open source cvxpy package (Diamond et al., 2014) to optimize the module weights. The objective is:
$\underset{w_{j}}{\operatorname{argmin}}\ E_{\mathcal{D}},\ \text{ s.t. }\sum_{j=1}^{N}w_{j}=1,\ w_{j}\geq 0$ (3)
Module Composition
After the optimization converges, we can use the learned weights to compose a single customized module for the specific task.
Inspired by the recent success of model merging in NLP and CV (Wortsman et al., 2022; Ilharco et al., 2022; Yu et al., 2024; Li et al., 2024; Yang et al., 2024a), we adopt a simple yet surprisingly effective method by weighted averaging the parameters of the selected modules:
$g_{\mathcal{D}}=\sum_{j=1}^{N}w_{j}^{*}g_{j},$ (4)
where $w_{j}^{*}$ represents the optimized weight for the $j$-th module in Equation 3. Here, the weights underscore the relevance of each selected module to the downstream task.
Table 1: Main results for 17 material property prediction tasks. The best MAE for each task is highlighted in bold and the second best result is underlined. Lower values indicate better performance. The results presented for each task are the average of five data splits, reported to three significant digits. For each method, the standard deviation of the test MAE across five random seeds is shown in parentheses. Additionally, the average rank and its standard deviation across the 17 datasets are provided to reflect the consistency of each method.
Datasets CGCNN MoE-(18) JMP-MT JMP-FT MoMa (Adapter) MoMa (Full)
Experimental Band Gap (eV) 0.471 (0.008) 0.374 (0.008) 0.377 (0.005) 0.358 (0.014) 0.359 (0.009) 0.305 (0.006)
Formation Enthalpy (eV/atom) 0.193 (0.015) 0.0949 (0.0016) 0.134 (0.001) 0.168 (0.007) 0.158 (0.009) 0.0839 (0.0013)
2D Dielectric Constant 2.90 (0.12) 2.29 (0.01) 2.25 (0.06) 2.35 (0.07) 2.31 (0.04) 1.89 (0.03)
2D Formation Energy (eV/atom) 0.169 (0.006) 0.106 (0.005) 0.140 (0.004) 0.125 (0.006) 0.112 (0.002) 0.0495 (0.0015)
Exfoliation Energy (meV/atom) 59.7 (1.5) 52.5 (0.8) 42.3 (0.5) 35.4 (2.0) 35.4 (0.9) 36.3 (0.2)
2D Band Gap (eV) 0.686 (0.034) 0.532 (0.008) 0.546 (0.020) 0.582 (0.018) 0.552 (0.014) 0.375 (0.006)
3D Poly Electronic 32.5 (1.1) 27.7 (0.1) 23.9 (0.2) 23.3 (0.3) 23.3 (0.2) 23.0 (0.1)
3D Band Gap (eV) 0.492 (0.008) 0.361 (0.003) 0.423 (0.004) 0.249 (0.001) 0.245 (0.002) 0.200 (0.001)
Refractive Index 0.0866 (0.0014) 0.0785 (0.0004) 0.0636 (0.0006) 0.0555 (0.0027) 0.0533 (0.0023) 0.0523 (0.0010)
Elastic Anisotropy 3.65 (0.11) 3.01 (0.03) 2.53 (0.26) 2.42 (0.36) 2.57 (0.61) 2.86 (0.28)
Electronic Dielectric Constant 0.168 (0.002) 0.157 (0.015) 0.137 (0.002) 0.108 (0.002) 0.106 (0.002) 0.0885 (0.0048)
Dielectric Constant 0.258 (0.008) 0.236 (0.002) 0.224 (0.004) 0.171 (0.002) 0.168 (0.002) 0.158 (0.002)
Phonons Mode Peak (cm-1) 0.127 (0.004) 0.0996 (0.0083) 0.0859 (0.0006) 0.0596 (0.0065) 0.0568 (0.0009) 0.0484 (0.0026)
Poisson Ratio 0.0326 (0.0001) 0.0292 (0.0001) 0.0297 (0.0003) 0.0221 (0.0004) 0.0220 (0.0003) 0.0204 (0.0002)
Poly Electronic 2.97 (0.10) 2.61 (0.13) 2.42 (0.03) 2.11 (0.04) 2.13 (0.03) 2.09 (0.03)
Poly Total 6.54 (0.24) 5.51 (0.04) 5.52 (0.03) 4.89 (0.06) 4.89 (0.04) 4.86 (0.07)
Piezoelectric Modulus 0.232 (0.004) 0.208 (0.003) 0.199 (0.002) 0.174 (0.004) 0.173 (0.003) 0.174 (0.001)
Average Rank 6.00 (0.00) 4.12 (1.17) 3.94 (0.97) 2.88 (1.27) 2.47 (0.94) 1.35 (0.86)
While alternative composition methods, such as mixture-of-experts, are feasible, they incur high memory overhead as MoMa Hub expands, limiting their practical deployment under computational constraints. By contrast, our weighted-average composition uses fewer resources while effectively integrating knowledge from all modules. In the full-module setting, every module shares the same architecture and pre-trained backbone with identical initializations, providing a grounded foundation for successful knowledge composition (Zhou et al., 2024).

### v3

# arXiv:2502.15483v3[cs.LG] 02 Mar 2026

@@ Abstract
Deep learning methods for material property prediction have been widely explored to advance materials discovery. However, the prevailing pre-train paradigm often fails to address the inherent diversity and disparity of material tasks. To overcome these challenges, we introduce MoMa, a simple Modular framework for Materials that first trains specialized modules across a wide range of tasks and then adaptively composes synergistic modules tailored to each downstream scenario. Evaluation across 17 datasets demonstrates the superiority of MoMa, with a substantial 14% average improvement over the strongest baseline. Few-shot and module scaling experiments further highlight MoMa’s potential for real-world applications. Pioneering a new paradigm of modular material learning, MoMa is open-sourced to foster broader community collaboration.
@@ 1 Introduction
Accurate and efficient material property prediction is critical for accelerating materials discovery. Key properties such as formation energy and band gap are fundamental in identifying stable and functional materials (Masood et al., 2023; Riebesell et al., 2025). While traditional approaches such as density functional theory offer high precision (Jain et al., 2016), their prohibitive computational cost limits their practicality for large-scale screening (Fiedler et al., 2022; Lan et al., 2023).
@@
Recently, deep learning methods have been developed to expedite traditional approaches (Xie and Grossman, 2018; Griesemer et al., 2023). Pre-trained force field models, in particular, have shown remarkable success in generalizing to a wide spectrum of material property prediction tasks (Shoghi et al., 2024; Rhodes et al., 2025; Wood et al., 2025), outperforming specialized models trained from scratch. These models are typically pre-trained on the potential energy surface (PES) data of materials (Barroso-Luque et al., 2024) and then fine-tuned for the target downstream task.
@@
Despite these advances, we identify two key challenges that undermine the effectiveness of current deep learning models for material property prediction: diversity and disparity.
Figure 1: Illustration of the diversity of material properties (top) and systems (down). Material tasks are also disparate, with different laws governing diverse properties and systems. These characteristics pose challenges for material property prediction models.
Figure 2: The modular learning scheme in MoMa trains and stores a broad spectrum of material tasks as modules, and adaptively composes them given a new material property prediction task.
@@
First, material tasks exhibit significant diversity (Fig. 2) which challenges the generalizability of existing models. For instance, prevailing force-field models are only trained on PES-derived properties (e.g., force, energy, and stress) mostly focusing on crystalline materials (Yang et al., 2024b; Barroso-Luque et al., 2024). However, material tasks span a much wider variety of systems (e.g., crystals, organic molecules) and properties (e.g., thermal stability, electronic behavior), making it difficult for methods trained on a limited set of data to generalize across the full spectrum of tasks.
@@
Second, the disparate nature of material tasks presents huge obstacles for jointly training a broad span of tasks in one model. Material systems vary significantly in atomic composition, bonding and structural periodicity, while their properties are governed by distinct physical laws. For example, mechanical strength in metals is primarily influenced by atomic bonding and crystal structure, whereas electronic properties like conductivity are determined by the material’s electronic structure. Consequently, training a single model across a wide range of tasks (Shoghi et al., 2024) may lead to knowledge conflicts, hindering the model’s ability to effectively adapt to downstream scenarios.
@@
Drawing inspiration from modular deep learning (Pfeiffer et al., 2023), we propose MoMa, a Modular framework for Material property prediction. Respecting the diversity challenge, MoMa trains multiple high-resource property prediction datasets into transferrable modules to support a wide-span of downstream tasks. In parallel, to address the disparity challenge, MoMa encapsulates each task within a specialized module during training to avoid interference. Furthermore, in adapting to each downstream task, MoMa adaptively integrates a synergistic combination of modules to mitigate knowledge conflicts. A high-level abstraction of MoMa is provided in Fig. 2.
@@
Specifically, MoMa comprises two major stages: (1) Module Training & Centralization. MoMa trains dedicated modules for a diverse range of material tasks, offering two versions: a full module for superior performance and a memory-efficient adapter module. These trained modules are centralized in MoMa Hub, a repository facilitating knowledge reuse while preserving proprietary data for privacy-aware material learning. (2) Adaptive Module Composition (AMC) & Fine-tuning. We devise AMC, a representation-driven, training-free module composition algorithm. Given a target task, AMC first estimates the performance of each module via kkNN label propagation in representation space. It then infers a weighted module composition by solving a convex optimization problem over a justified proxy error. The composed module is then fine-tuned on the downstream data for improved adaptation. Together, the two stages offers a flexible and scalable solution to achieve effective modular learning for material property prediction.
@@
Empirical results across 17 downstream tasks showcase the superiority of MoMa, outperforming all baselines in 16/17 tasks, with an average improvement of 14% compared to the best non-modular baseline. In few-shot settings, which are common in materials science, MoMa achieves even larger performance gains to the conventional pre-train then fine-tune paradigm. Additionally, MoMa shows improved average improvements as we scale the number of modules in the MoMa Hub, and the AMC-optimized weights provide valuable insights into relationships between material properties. The code and trained modules of MoMa are open-sourced, and we envision MoMa becoming a pivotal platform for the modularization and distribution of materials knowledge, fostering deeper community engagement to accelerate materials discovery.
2 Related Work
2.1 Material Property Prediction with Deep Learning
Deep learning methods have been widely adopted for predicting material properties (De Breuck et al., 2021). The seminal CGCNN model (Xie and Grossman, 2018) represents crystalline materials with multi-edge graphs and applies graph neural networks for representation learning. Subsequent work (Choudhary and DeCost, 2021; Das et al., 2023; Yan et al., 2024; Taniai et al., 2024) has focused on improving neural network architectures to better model the inductive biases of crystals.
Another line of work develops pre-training strategies for materials (Jha et al., 2019; Magar et al., 2022; Wang et al., 2025). Recently, a series of large force field models (Merchant et al., 2023; Batatia et al., 2023; Neumann et al., 2024) are trained on massive Potential Energy Surface data (Barroso-Luque et al., 2024) and achieve remarkable accuracy in material tasks (e.g. thermal stability prediction (Riebesell et al., 2025)). Notably, the JMP model (Shoghi et al., 2024), trained across multiple domains (small molecules, catalysts, etc.), performs impressively when fine-tuned on both molecular and crystalline tasks.
Extending beyond these methods, MoMa offers a modular strategy to centralize diverse material knowledge into modules and adaptively compose them, yielding superior downstream performance.
2.2 Modular Deep Learning
Modular deep learning (Pfeiffer et al., 2023; Xiao et al., 2024) represents a promising paradigm where parameterized modules are composed, selected, and aggregated for function specialization and reuse. Notable examples of modular networks include mixture-of-experts (Jacobs et al., 1991; Shazeer et al., 2016), adapters (Houlsby et al., 2019) and LoRA (Hu et al., 2021). Recently, we have seen increasing applications of modular methods across domains such as NLP (Pfeiffer et al., 2020; Huang et al., 2024; Tan et al., 2024) and CV (Puigcerver et al., 2020; Pham et al., 2024), where its strengths in flexibility and minimizing negative interference have been demonstrated.
An important aspect of modular learning is how modules are weighted prior to composition. Previous adaptive module composition approaches can be broadly grouped into (1) search-based methods that iteratively optimize weights based on downstream predictive performance after composition (Huang et al., 2024; Akiba et al., 2025), and (2) router-based methods that learn composition weights via an additional routing network (Muqeeth et al., 2023; Lu et al., 2024). Crucially, both paradigms rely on the downstream prediction error of the composed model to guide weight allocation. However, this dependence is problematic in material settings: high task disparity makes the error signals (from arbitrary module mixtures) noisy and unstable for search-based methods, while data-scarcity provides insufficient supervision for router learning. Additionally, loading all material modules during router training becomes prohibitively costly as the number of module scales.
In the context of material property prediction, modular learning remains largely under-explored. The most related work is the router-based mixture-of-experts method MoE-(18) (Chang et al., 2022), which loads all available modules and learns a routing network for embedding aggregation.
@@ 3 Proposed Framework: MoMa
MoMa is a simple modular framework targeting the diversity and disparity of material property prediction tasks. MoMa involves two major stages. In the first stage (Section 3.1), we train and centralize modules for a diverse range of material systems and properties into MoMa Hub. In the second stage (Section 3.2), we devise a representation-driven, training-free algorithm to adaptively select and compose MoMa hub modules for a target task, and then fine-tune the composed model. A visual overview of MoMa is shown in Figure 3.
Figure 3: The MoMa framework. (a) During the Module Training & Centralization stage (Section 3.1), MoMa trains full and adapter modules for a wide spectrum of material tasks, constituting the MoMa Hub; (b) The Adaptive Module Composition (AMC) & Fine-tuning stage (Section 3.2) leverages the modules in MoMa Hub to compose a tailored module for each downstream task. The AMC algorithm comprises three steps: 1. Prediction Estimation; 2. Weight Optimization; 3. Module Composition. The composed module is further fine-tuned on the task for better adaptation.
@@ 3.1 Module Training & Centralization
We provide two parametrizations for the MoMa modules: the full module and the adapter module. For the full module, we directly treat each fully fine-tuned model backbone as a standalone module. The adapter module, in contrast, serves as a parameter-efficient alternative where adapter layers (Houlsby et al., 2019) are inserted between each layer of the backbone. The adapters are updated and the rest of the backbone is frozen. All adapters trained for a given task are collectively treated as one module. This implementation trade-offs the downstream performance for a much lower GPU memory cost during training, making it especially suitable for compute-constrained settings. When training converges, all module parameters are stored into a centralized repository $\mathcal{H}$ termed MoMa Hub, formally:
$\mathcal{H}=\{g_{1},g_{2},\dots,g_{N}\},\quad g_{i}=\begin{cases}\theta_{f}^{i}&\text{(full module)}\\ \Delta_{f}^{i}&\text{(adapter module)}\end{cases}$
where $\theta_{f}^{i}$ and $\Delta_{f}^{i}$ denote the full and adapter module parameters for the $i^{\text{th}}$ task and encoder $f$.
@@
To support a wide array of downstream tasks, MoMa Hub needs to include modules trained on diverse material systems and properties. Currently, MoMa Hub encompasses 18 material property prediction tasks selected from the Matminer datasets (Ward et al., 2018) with over 10000 data points. These tasks span across a large range of material properties, including thermal properties (e.g. formation energy), electronic properties (e.g. band gap), mechanical properties (e.g. shear modulus), etc. For more details, please refer to Table 5. Note that MoMa is designed to be task-agnostic and may readily support a larger spectrum of tasks in the future.
An important benefit of the modular design of MoMa Hub is that it preserves proprietary data, which is prevalent in the field of materials, enabling privacy-aware contribution of new modules. Therefore, MoMa could serve as an open platform for the modularization of materials knowledge.
@@ 3.2 Adaptive Module Composition & Fine-tuning
Given a labeled material property prediction dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{M}$, the goal of the second stage is to customize a task-specific model by composing modules from MoMa Hub. Due to the diversity and disparity of material tasks, blindly combining modules often leads to suboptimal performance. The composition must be adaptive, composing only the most synergistic modules for each task. Furthermore, given the vast and expanding scale of the Hub, the method must be data-driven and efficient, avoiding reliance on human expertise or prohibitively expensive exhaustive search.
However, satisfying these requirements is non-trivial for existing adaptive weighting paradigms. As discussed in Section 2.2, both search-based and router-based methods rely on downstream prediction error derived from composed module as the supervision signal. In our setting, this signal is less reliable: the high disparity of modules in inputs (e.g. crystals vs. molecules) and targets (e.g. energies vs. band gaps) induces highly heterogeneous representation spaces. Hence module mixtures yield unstable representations and uninformative error signals, resulting in a noisy optimization landscape that hampers search-based methods. Moreover, the scarcity of downstream data makes router training difficult and prone to overfitting.
To address these limitations, we devise the Adaptive Module Composition (AMC) algorithm. Instead of relying on prediction error supervision, AMC adopts a representation-driven and training-free strategy. Specifically, it first estimates per-module performance via kkNN in the representation space, and then solves for optimal composition weights by minimizing an ensemble proxy error via convex optimization. This allows AMC to efficiently identify synergistic compositions without iterative search or extra trainable parameters. We now introduce AMC in detail. An analytical figure of AMC is provided in Fig. 5, with its formal formulation in Algorithm 1.
Figure 4: An analytical decomposition of AMC. Blue arrows: per-module kkNN prediction estimation in representation space on target task. Black arrows: convex optimization of ensemble proxy error to obtain composition weights. Orange arrows: weight-space module composition to construct the final composed module.
Figure 5: Scatter plot showing the relationship between kkNN proxy MAE and post–fine-tuning MAE of standalone MoMa Hub modules on three representative tasks. Colored lines are linear fits. We observe a clear positive correlation with Pearson’s $r>0.6$.
Representation-driven Prediction Estimation
AMC begins by estimating the affinity of each module to the downstream task. To bypass the unstable optimization landscape of arbitrary module mixtures, we first evaluate the intrinsic representation quality of each module individually. We posit that a task-aligned module should map materials with similar properties to adjacent points in the embedding space.
Formally, for each module $g_{j}\in\mathcal{H}$, we encode the training data $\mathcal{D}$ into representations $\mathcal{X}^{j}=\{\mathbf{x}_{1}^{j},\ldots,\mathbf{x}_{M}^{j}\}$. We then perform leave-one-out kkNN label propagation (Iscen et al., 2019) to obtain a prediction $\hat{y}_{i}^{\,j}$ for each instance:
$\hat{y}_{i}^{\,j}=\sum_{k\in\mathcal{N}_{i}}\frac{f_{d}\!\left(\mathbf{x}_{i}^{\,j},\,\mathbf{x}_{k}^{\,j}\right)}{Z_{i}^{\,j}}\,y_{k},\quad Z_{i}^{\,j}=\sum_{k\in\mathcal{N}_{i}}f_{d}\!\left(\mathbf{x}_{i}^{\,j},\,\mathbf{x}_{k}^{\,j}\right).$ (1)
@@
where $\mathcal{N}_{i}$ denotes the indices of the $K$ nearest neighbors of $\mathbf{x}_{i}^{j}$ within $\mathcal{X}^{j}$, and $f_{d}$ is the exponential cosine similarity function.
We choose kkNN as the estimator because it directly probes the local geometry of the representation space without introducing learnable parameters. This strictly aligns with our training-free design principle and ensures robustness against overfitting on data-scarce tasks.
Training-free Module Weight Optimization
With the module-wise performance estimates $\{\hat{\mathbf{y}}^{j}\}_{j=1}^{N}$ from the representation space, our goal is to identify an optimal weight vector $\mathbf{w}\in\mathbb{R}^{N}$ (where $w_{j}$ denotes the weight of module $j$) to compose these modules. While the ideal objective is to minimize the validation error of the fine-tuned model, searching this space directly is computationally infeasible due to combinatorial explosion. Instead, inspired by ensemble learning (Zhou et al., 2002; Zhou, 2016), we propose to use the prediction error of the weighted ensemble (prior to fine-tuning) as a proxy error to guide weight optimization.
Specifically, we formulate the composition prediction as a weighted sum of the individual module estimations. The proxy error $E_{\mathcal{D}}$ is defined as the mean squared error between the ensemble prediction and the ground truth labels on the training set:
$E_{\mathcal{D}}(\mathbf{w})=\frac{1}{M}\Big\|\sum_{j=1}^{N}w_{j}\hat{\mathbf{y}}^{j}-\mathbf{y}\Big\|_{2}^{2}.$ (2)
We can further cast Eq. 2 to a constrained optimization problem:
$\underset{\mathbf{w}}{\operatorname{argmin}}\ E_{\mathcal{D}}(\mathbf{w}),\quad\text{ s.t. }\ \sum_{j=1}^{N}w_{j}=1,\ w_{j}\geq 0.$ (3)
Since the objective is convex and the constraints define a convex feasible set, the problem admits a global optimum that can be reliably obtained by standard solvers. Moreover, this weight selection is training-free since it introduces no additional learnable parameters and requires no gradient-based updates or hyperparameter tuning beyond the optimizer settings.
Justification for Using the Proxy Error
A central premise of AMC is that the kkNN-based proxy error (Eq. 2 is a reliable indicator of the final model performance. Theoretically, we provide a formal risk analysis in Appendix B to show that, under reasonable assumptions, minimizing this proxy error bounds the risk of the subsequently fine-tuned model. Empirically, when measured in MAE to align with downstream metrics, we observe a strong Pearson correlation (>0.6>0.6) between the per-module proxy errors and their post-fine-tuning performance (see Fig. 5 and Section D.1 for detailed discussion). This indicates that the proxy is a reliable predictor of final performance and supports its use for weight optimization.
Weight-space Module Composition
Once the optimal weight vector $\mathbf{w}^{*}$ is obtained, we compose a single customized module $g_{\mathcal{D}}$ for the target task. Inspired by recent advances in model merging (Wortsman et al., 2022; Ilharco et al., 2022; Yu et al., 2024; Yang et al., 2024a), we merge the modules in their weight space: $g_{\mathcal{D}}=\sum_{j=1}^{N}w_{j}^{*}g_{j}.$
The validity of this averaging is supported by linear mode connectivity (Frankle et al., 2020; Zhou et al., 2023; 2024). Since all modules originate from a common pre-trained initialization, their parameters remain structurally compatible despite task-specific divergence. This ensures that the composed module serves as a stable and well-conditioned initialization for downstream fine-tuning.